Zimmer Company sold the following two machines in 2020:
Machine A Machine B
Cost $76,000 $80,000
Purchase date July 1, 2016 January 1, 2017
Useful life 8 years 5 years
Salvage value $4,000 $4,000
Depreciation method Straight-line Double-declining-balance
Date sold July 1, 2020 August 1, 2020
Sales price $35,000 $16,000
Instructions
1. Pass the necessary journal entries to record purchase of the machine A and
Machine B on 1 July 2016 and 1 January, 2017 respectively.
2. Calculate depreciation of the machine A and machine B for the year 2016, 2017,
2018, 2019. .
3. Journalize all entries required to update depreciation and record the sales of the
two assets in 2020.
Answer #1

1.

Date General Journal Debit Credit
01-07-16 Machine A            76,000
Cash          76,000
(being purchase of Machinary for cash)
01-01-17 Machine B          80,000
Cash          80,000
(being purchase of Machinary for cash)

2.

Compuation of Depreciation
Particulars Machine A Machine B
Purchase Date 01-07-16 01-01-217
Cost Of Machine (a)                       76,000                             80,000
Salvage Value (b)                         4,000                               4,000
Useful life (c ) 8 Years 5 Years
Depreciation rate (d) 12% 40%
(((a-b)/c)/a) ((a/c)*2)
Asset Used in 2016 (e) 6 months                                     -  
Depreciaton for 2016 (f)                         4,500                                     -  
(a*d)*e/12                                     -  
Asset used in 2017 (g) 1 Year 1Year
Depreciation for 2017 (h)                         9,000                             32,000
(a*d) (a*d)
Written down value in 2017 (i=a-f-h)                       62,500                             48,000
Depreciation for 2018 (j)                         9,000                             19,200
(a*d) (i*d)
Written down value in 2018 (k=i-j)                       53,500                             28,800
Depreciation for 2019 (l)                         9,000                             11,520
(a*d) (k*d)
Written down value in 2019 (m=k-l))                       44,500                             17,280
Asset used in 2020 (n) 6 months 8 months
Depreciation in 2020 (o)                         4,500                               6,912
(a*d) (m*d)
Written down value on sale date (p=m-o)                       40,000                             10,368
(as on July 1 2020) (as on August 1 2020)
Sale Amount (q)                       35,000                             16,000
Profit / (Loss) on Sale (r=q-p)                       (5,000)                               5,632

3.

Date General Journal Debit Credit
01-07-20 Depreciation            4,500
Machine A            4,500
(being depreciation booked for till july 2020)
01-07-20 Cash          35,000
Loss on sale of Fixed Asset            5,000
Machine A            5,000
(being sale of Machine A)
01-08-20 Depreciation            6,912
Machine B            6,912
(being depreciation booked for till August 2020)
01-08-20 Cash          16,000
Machine B          10,368
Profit on Sale of Fixed Asset            5,632
(being sale of Machine B)

I assumed that Asset shown in balance sheet as written down value instead of showing accumulated depreciation in balance sheet